Course Evaluation

Course Evaluation – p.1/7

What is the process?

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Instructor maintains a binder of course material (lectures, quizzes, midterms, final project, ...) during the semester.

Students fill out SPOT forms.

At the end of the semester, instructor fills out the Course Assessment form.

At the beginning of the summer term, a team of faculty members evaluate each course based on the above evidence and suggest changes.

Every six years, the entire undergraduate program is evaluated by ABET for accreditation.

Every six years, the university evaluates both undergraduate and graduate programs to see if the programs should be continued (QER).

Course Evaluation – p.2/7

Is student input important?

Absolutely Student input is useful to make adjustments not only to this course but also to its pre-requisites. Students are good at spotting course deficiencies. Students are good at spotting course enhancements.

Absolutely Student input is useful to make adjustments not only to this course but also to its pre-requisites. Students are good at spotting course deficiencies. Students are good at spotting course enhancements.

Absolutely Student input is useful to make adjustments not only to this course but also to its pre-requisites. Students are good at spotting course deficiencies. Students are good at spotting course enhancements.

What are the specifics of this course?

Other Aspects

What happens to the evaluation forms?

Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project Development of process flowsheets Mass and energy balances of entire flowsheets Equipment sizing and costing Optimization The use of process simulators (CHEMCAD) The use of computer tools (EXCEL, MATLAB, CAPCOST) “Putting it all together” in a design project The use of Blackboard for the entire course The use of software in class Movie clips Instructor availability The use of Blackboard for the entire course The use of software in class Movie clips Instructor availability The use of Blackboard for the entire course The use of software in class Movie clips Instructor availability The use of Blackboard for the entire course The use of software in class Movie clips Instructor availability Evaluation forms are sent to the university to be processed.

If over 90% of the students indicate that the instructor is “excellent” in the last question of section D, the instructor gets a letter of commendation from the provost.

If over 20% of the students indicate that the instructor is “poor” in the last question of section D, the instructor gets a phone call from the provost.

Evaluation forms are sent to the university to be processed.

If over 90% of the students indicate that the instructor is “excellent” in the last question of section D, the instructor gets a letter of commendation from the provost.

If over 20% of the students indicate that the instructor is “poor” in the last question of section D, the instructor gets a phone call from the provost.

Evaluation forms are sent to the university to be processed.

If over 90% of the students indicate that the instructor is “excellent” in the last question of section D, the instructor gets a letter of commendation from the provost.

If over 20% of the students indicate that the instructor is “poor” in the last question of section D, the instructor gets a phone call from the provost.

What else can students do?

If they think that the instructor did an outstanding job, they can nominate him/her for a teaching award.

If they think that the instructor did a lousy job, they can complain (in writing) to the department chair.

What else can students do?

If they think that the instructor did an outstanding job, they can nominate him/her for a teaching award.

If they think that the instructor did a lousy job, they can complain (in writing) to the department chair.